# references.bib Weak Ties Explain Open Source Innovation

### v1

@@ III-B2 Interactions—Commits, Issues, Stars
Consequently, we construct three directed topological project-to-project networks, one for each of the three interaction types, to represent possible flow of knowledge across projects. We record an edge if and only if a core developer of project interacted with project within one year prior to their initial commit to project , as illustrated in Figure 2. The weight of each edge represents the number of core developers from project who engaged with project in a specific type of interaction. The one-year window models that recent interactions may be more relevant.
III-B3 Validation
To further validate that interaction via commits, issues, and stars is a good proxy for ties of decreasing strength, we empirically compare the transitivity of the three networks. First, consider the illustration in Figure 3: if node is connected to and , a triad is formed. In social network theory, if ’s connections to and are both strong ties, the triad is unstable and will likely evolve into a triangle where , , and are all connected. This evolution occurs because and have more opportunities to interact and increase their mutual familiarity due to their strong ties with . Thus, networks characterized by strong ties typically exhibit a greater number of triangles, as opposed to triads, compared to those with weak ties [granovetter1973strength]. Formally, this can be computed using the measurement of transitivity, defined as . A higher transitivity value (number of triangles relative to triads) indicates a network of stronger ties.
Similarly, we compare the transitivity values of our three networks, summarized in Table I.111Since network transitivity is usually defined for undirected networks, we first convert our three networks (which are all directed) to their undirected form. That is, there is an undirected edge between nodes and if there is either an edge from to or from to . In addition, the reported number of nodes excludes isolates (i.e., nodes without edges) from each network, and the number of edges is calculated in the undirected version of each network. Overall, we observe approximately an order of magnitude () difference in transitivity values between each pair of networks. Specifically, the commit network displays the highest levels of transitivity, followed by the issue network, while the star network exhibits the lowest level. Thus, these findings are consistent with our theoretical understanding of tie strength.
| Interaction | #Nodes | #Edges | Transitivity () |
| Commits | |||
| Issues | |||
| Stars |
III-C Network Measures
Next we describe the variables we computed to capture a project’s position in the different networks.
III-C1 Using Out-Degree Centrality to Measure the Amount of Interaction and Accessible Knowledge
For every project, we compute its out-degree centrality within each of the three networks (commits, issues, and stars), defined as the total sum of the weights of all its out-edges. These out-edges represent the number of actions initiated by the core developers of the project towards other projects in the past, and are a proxy for the amount of knowledge from other projects available to the core developers of the focal project; this knowledge may end up influencing the design of the focal project.
This step resulted in computing the three variables listed in Table IV under Degree Variables (original).
III-C2 Using A Node2Vec Model to Measure the Diversity of Accessible Knowledge
Next, we compute a proxy for the informational diversity of projects to which a focal project is connected within each interaction network. This diversity metric reflects the extent to which the focal project can access knowledge from a variety of sources. While conventional metrics like network constraint [burt1997note] and betweenness centrality [brandes2001faster] are commonly used to measure network diversity, they often do not scale well and can be computationally intensive, especially when dealing with large networks such as those in our study. Moreover, these metrics typically consider both the in-degree and out-degree edges of the focal node, while our study relies solely on the out-degree to represent the possible flow of knowledge from other projects to the focal one.
@@ III-C2 Using A Node2Vec Model to Measure the Diversity of Accessible Knowledge
As a result, we use a Node2Vec model [grover2016node2vec], a classical approach from the machine learning domain, to generate vector representations (embeddings) for each node in the network based on its topological position. Initially, we generate sequences of random nodes from the graph using a random walk process: Starting from a random node in the network, we sample the next node by randomly selecting another node in the network that is adjacent to the current one. The likelihood of selecting each node is proportional to the edge weight between the candidate next node and the current node. This process continues until we have selected enough nodes to complete the walk (in our study, each walk comprises 20 nodes). We repeat this process to generate multiple walks (or sequences of nodes).
Subsequently, we use a skip-gram model [mikolov2013efficient], commonly used for generating embeddings for words in natural language, on the generated walks to obtain embeddings for each node. Nodes that frequently appear together in the same walk, representing sets of projects that are related in some informational sense, tend to have vector representations that are closely situated in the embedding space.
Next, for each project in the network, we examine all other projects that receive a directed edge from the focal project (denoted as set ). Our knowledge diversity index is the pairwise distance between any two projects in :
where and are projects in , and and are their vector representations from the Node2Vec model. The distance between and is the negation of the cosine similarity of the two vectors.
We calculate the diversity index for each node in each network. However, the index is undefined for projects with an out-degree centrality less than two. Therefore, such projects will be excluded from our regression analysis when considering diversity metrics as independent variables.
This step resulted in computing the three variables listed in Table IV under Diversity Variables (original).
III-C3 Validation & Exploratory Analysis
One can expect that the three degree variables and the three diversity variables are correlated to some extent, since they are computed in the same way (albeit on different networks). To prevent issues of multicollinearity in our subsequent regression models, which can both complicate the interpretation of the estimated coefficients and reduce the statistical power of our model [daoud2017multicollinearity], we first use Principal Component Analysis (PCA) [wold1987principal] to decompose the correlated variables into orthogonal principal components.
@@ III-C3 Validation & Exploratory Analysis
Since they are designed to capture different concepts, we run PCA separately for the out-degree centrality variables across the three networks (referred to as the degree variables) and the structural diversity measurements across the three networks (referred to as the diversity variables).222Recall that for the diversity variables, the dataset (hence also PCA) is restricted to projects with an out-degree of at least two in all three networks, since otherwise the variables are undefined. As standard, before running PCA we first log-transformed all input variables and scaled them to a mean of zero and a standard deviation of one.
@@
The PCA resulted in three principal components for the degree variables and another three for the diversity variables, with the proportion of variance explained by each component listed in Table II. Inspecting the table, we observe that the first two PCs cumulatively explain over 80% of the variance in both groups, thus we decided to retain only two principal components each for the degree and diversity variables. That is, we will use these PCs (listed under Degree Variables (post-PCA) and Diversity Variables (post-PCA) in Table IV) instead of the original variables in our regression models below.
| Out-deg. centrality | Diversity index | |||||
| PC1 | PC2 | PC3 | PC1 | PC2 | PC3 | |
| Variance Explained | 0.64 | 0.22 | 0.14 | 0.51 | 0.29 | 0.19 |
| Cumulative Variance | 0.64 | 0.86 | 1.00 | 0.51 | 0.81 | 1.00 |
| Out-deg. centrality | Diversity index | |||||
| PC1 | PC2 | PC3 | PC1 | PC2 | PC3 | |
| 0.60 | 0.45 | 0.67 | 0.63 | 0.36 | 0.69 | |
| 0.61 | 0.28 | 0.74 | 0.64 | 0.24 | 0.72 | |
| 0.52 | 0.85 | 0.11 | 0.43 | 0.90 | 0.08 | |
@@
To interpret the two PCs we turn to Table III, which lists the loadings of the principal components onto the original variables.333Principal components are linear combinations of the original variables, with the loadings indicating the contributing coefficients. For example, the first PC for the degree variables, , uses the coefficients in the first column in Table III: . Inspecting the table we make the following observations. First, the loadings onto the first principal component (PC1) are relatively consistent across networks for both the degree and the diversity variables. Thus, we interpret PC1 to represent the average degree (or diversity of ties) of the project across the three networks. For example, drawing on our theoretical framework (Section II), a project for which is high is expected to have many ties (on average, across the three networks), i.e., its core developers should have more sources to draw inspiration from. Similarly, the ties of a project for which is high are expected to span more of the knowledge embedding space we used to estimate tie diversity, i.e., its core developers could have access to more varied sources to draw inspiration from.
@@
Second, the loadings onto the second principal component (PC2) exhibit a descending trend across networks, again similarly for both the degree and the diversity variables. Notably, the highest loading on the degree (or diversity) comes from the star network, followed by the issue network, while the lowest comes from the commit network. Thus, we interpret PC2 to represent the strength of network ties in the degree (or diversity of ties) metric, i.e., the strength of weak ties. A project for which is high is expected to get more of its connectivity through the star network. Analogously, the diversity of knowledge accessible to core developers in a project for which is high can be attributed relatively more to the star network (weak ties) compared to the commit or issue networks (stronger ties).
@@
It may seem counterintuitive to reason about PC1 and PC2 jointly, especially in a regression modeling framework like we apply below. What does it mean to vary PC2 from low to high, for instance, while holding PC1 fixed? Can PC1 and PC2 even vary independently? They can in theory, because they are orthogonal by construction, but are the combinations of low-high values of PC1 and PC2 even observable in the real-world data? After all, regression is an interpolation mechanism. We illustrate this space of interpretations of PC1 and PC2 for the diversity variables with the artificial examples in Figure 4 (see our replication package for real-world examples corresponding to each quadrant). In the top left quadrant (PC1 and PC2 both high), the project core developers can tap into a diverse knowledge space via either strong or weak ties. In the top right (PC1 high and PC2 low), the overall diversity of information accessible is just as high, but it’s the strong ties that are the source of it. In the bottom left (PC1 low and PC2 high), the overall diversity of information available is lower, but whatever diversity there is, it can be attributed mostly to the weak ties. Finally, in the bottom right (PC1 and PC2 both low), weak ties contribute little to an otherwise low diversity of information.
@@
Going back to our hypotheses (Section II), we expect the left quadrants, where weak ties are strong, to be most associated with innovativeness. As an anecdote, consider the example in Figure 1 on the first page. The GeoNotebook project is a Jupyter notebook-based environment for interactive visualization and analysis of geographic data. Interestingly, GeoNotebook’s core developers have previously starred (without otherwise contributing to) many GitHub projects spanning a variety of seemingly related topics, including a collection of free APIs, example data science notebooks, reproducible workflows for Jupyter notebooks, visualization tools, time series analysis tools, geographic data analysis tools, and infrastructure tools for interactive packet manipulation. One can expect that this diverse space of ideas provided some inspiration, knowingly or unknowingly, for the design and implementation of GeoNotebook. Our analysis below tests to what extent there is evidence supporting this mechanism at scale, beyond this anecdote.
@@
| Innovativeness | The extent to which a project imports software packages that were typically not imported together in the past, computed based on all the packages imported until the end of year 2021. |
@@
| The total number of core developers in the project, as computed before the end of year 2021. | |
| The number of software packages that the project imported before the end of year 2021. | |
III-D Project Innovativeness Measure
fang2024novelty introduced a measure of a software project’s novelty, or innovativeness, as a function of the combinations of packages the project imports. For example, while certain libraries, such as numpy and tensorflow, are frequently used together, others, like numpy and requests, are less commonly combined. The measure estimates how atypical combining two packages is, and by extension how atypical is the overall set of packages being used as dependencies in a project, by comparison to what could be expected by random chance. Projects that import more atypically paired packages are deemed more innovative as they reuse packages in nontraditional ways.
[…]
@@ III-D Project Innovativeness Measure
Next, we determine the atypicality of a combination of two packages by computing the negation of the cosine similarity of their embeddings. Finally, we define our measure of project innovativeness as the average pairwise atypicality of all pairs of packages imported by a project.444Recall that we can only compute the atypicality score for projects that import at least two packages. Other packages are excluded from the analysis.
[…]
III-E Regression Modeling Considerations
@@
To test our three hypotheses, we use fixed-effects linear regression (one data point per project), where the response variable is our measure of innovativeness, and our main explanatory variables are the degree and diversity principal components discussed above: captures H (volume of interactions the core developers had with other developers and projects in the previous year), captures H (diversity of information accessible), and captures H (strength of weak ties). The remaining principal component, , does not directly map to one of our hypotheses, but can help distinguish whether it’s the volume of weak ties, or their strength, that explains more variance in our outcome measure.
@@
In addition, we control for several important confounding variables, as described in Table IV: the repository owner’s social standing (), the project team size (), the complexity of the codebase in terms of number of packages imported (), whether the project is owned by an organizational account (Org_owned), and the project age (, modeled as a categorical variable). For model fit and diagnostics we followed standard practice, e.g., we first log-transformed the variables with skewed distributions to reduce heteroskedasticity [kaufman2013heteroskedasticity]; see replication package.
IV Results
@@
Next we present the main results from the regression analysis formally testing our hypotheses, summarized in Table V, as well as a series of robustness checks.
@@ IV-A Main Regression Results
Model I (all projects) and Model II (projects with a minimum out-degree centrality of two across all three networks, such that the diversity variables are well defined) shed light on the impact of degree variables. Model I reveals a significant positive effect from average degree centrality, whereas Model II shows a negative effect. Nevertheless, effect sizes for both models are relatively small. For instance, in Model I, transitioning a project’s from the 25th percentile (-1.08) to the 75th percentile (0.54) results in a mere 0.002 increase in project innovativeness. Considering that the 25th percentile of a project’s innovativeness score is -0.33, and the 75th percentile is -0.12 in our sample, the change in project innovativeness due to variations in average degree centrality is practically negligible. Therefore, H is not supported.
@@
In Model III, we observe a positive association between the average diversity () of ties originating from the focal project and project innovativeness. Moreover, this effect is notably stronger—nearly seven times greater—compared to that of average degree in Models I and II. Transitioning a project from the 25th percentile (-0.89) to the 75th percentile (0.86) in terms of corresponds to an increase of approximately 0.012 in project innovativeness, representing roughly a 4% change in the distribution (e.g., a shift from the median project in terms of novelty to the 54th percentile). We conclude that H is supported.
@@
Additionally, we observe a significant effect for the variable. This suggests that, when controlling for the average diversity across the three networks, higher diversity in the “weak ties network” (e.g., star network), or stronger weak ties, corresponds to increased project innovativeness. In practical terms, the effect size for the variable is comparable, albeit somewhat smaller, to that of the variable above. We take this as evidence that H is supported.
@@
When incorporating both degree and diversity variables in the same regression model, as in Model IV, the positive effects of and remain significant. This suggests a consistent effect stemming from tie diversity and strength of weak ties. Nevertheless, we also note a significant negative effect for and after adjusting for project diversity. These negative effects suggest that, while holding tie diversity constant, higher out-degree centrality and more ties in the “weak network” are associated with decreased project innovativeness. This observation could be attributed to the increased constraints within a local community associated with higher out-degree, as discussed by Burt [burt2000network]. Such constraints can lead to a reduced inclination for combining packages of diverse functions within the project, consequently contributing to the reduced novelty of the project. More research is needed to further test these effects.
IV-B Robustness Checks
@@
In addition to the regressions above, we test how robust our conclusions about the effects of and are to differences in operational decisions.
IV-B1 Does The Time of Measurement Matter?
@@
We segmented our dataset by project inception year and conducted regression analyses with the specification of Model III in Table V on projects initiated in each respective year. Figure 5(a) shows the resulting estimated coefficients for the two diversity variables. As illustrated in the graph, the positive relationship between the average diversity of knowledge access across the three networks (represented by the blue line) remains consistent over the years, with the average estimated coefficient being statistically significantly above zero in most years. The estimated coefficient for appears to also increase in recent years, though this difference is not statistically significant.
@@
The positive relationship between diversity from weak ties and project innovativeness also appears generally persistent (the red line in Figure 5(a)). The estimated coefficient is above zero in nine out of the ten years observed, and the statistically insignificant coefficient (i.e., the lower bound of the 95% confidence interval is below zero) is likely attributable to the limited number of projects in each annual cohort (i.e., fewer than 4,000 projects in each year before 2016).
IV-B2 Does The Project Size Matter?
@@
Similarly, we split projects into subsets based on the number of core developers and assessed the impact of the diversity variables in each cohort. As shown in Figure 5(b), the effect of remains consistent across projects of varying team sizes, confirming the robustness of our findings. It is noteworthy that the effect of becomes insignificant for projects with four or more core developers. This may be attributed to the limited number of project samples within these categories.
IV-B3 Does The Project Ownership Matter?
@@
Lastly, we re-estimated the effects of the diversity variables separately for projects owned by organizational and individual accounts. Our analysis (Figure 5(c)) shows that both average diversity across networks and diversity from weak ties have significantly positive effects for both categories. Furthermore, the difference in effect size between organization-owned and individual-owned projects is not statistically significant.
In conclusion, the positive relationships we observed between average knowledge diversity and project innovativeness, as well as between diversity from weak ties and project innovativeness, are generally robust.
V Discussion

### v2

@@ 3.3. Network Construction
Each of the three actions we consider (making commits, participating in issue discussions, and starring) can be thought of as a person–to–project tie. Aggregating across all developers and projects in our sample leads to three person–to–project bipartite networks, one for each action type. However, since our analysis is at the project level, we construct three directed project–to–project networks by way of projection, to represent possible flow of knowledge across projects. As discussed previously, we consider only actions initiated by core developers (e.g., which other repositories they committed to or starred in the past). In addition, we restrict the scope to actions in the recent past (e.g., within the last year, we validate with alternative period length in the appendix), as we expect older actions to be less relevant. This time window is developer-dependent. That is, we record an edge if and only if a core developer of project interacted with project within one year prior to their initial commit to project , as illustrated in Figure 2. The weight of each edge represents the number of core developers from project who engaged with project via a specific type of interaction.
As expected, the structural properties of the three networks support our reasoning for why commits, issues, and stars, respectively, capture ties of decreasing strength. Comparing the transitivity222In network science, transitivity is a measure of the ratio of triangles to triads. values of our three networks, we observe approximately an order of magnitude () difference between each pair of networks. Specifically, the commit network displays the highest levels of transitivity, followed by the issue network, while the star network exhibits the lowest level; see Supplementary Materials for details.
3.4. Network Measures
Next, we compute variables related to the network position of each project to represent the amount and diversity of knowledge that may be accessible to its core developers through each network.
First, we use the number of connected projects as a proxy for the amount of knowledge accessible to a given project. More formally, for every project, we compute its out-degree centrality within each of the three networks, defined as the total sum of the weights of all its out-edges. This step resulted in computing the three variables listed in Table 2 under Degree Variables (original).
@@ 3.4. Network Measures
Second, we use the Node2Vec graph embedding algorithm (Grover and Leskovec, 2016) to generate vector representations (embeddings) for each node in the network based on their topological position. The process begins by generating sequences of random nodes from the graph using random walks. Starting from a random node in the network, we sample the next node by randomly selecting a direct neighbor. The likelihood of selecting each node is proportional to the edge weight between the candidate next node and the current node. This process continues until we have selected enough nodes to complete the walk (in our study, each walk comprises 20 nodes). We repeat this process to generate multiple walks (or sequences of nodes). Subsequently, we use a skip-gram model (Mikolov et al., 2013a), commonly used to generate embeddings for words in natural language, on the generated walks to learn embeddings for each node.
Next, for each project in the network, we examine all other projects that receive a directed edge from the focal project (denoted as set ). Our knowledge diversity index333The index is undefined for projects with out-degree centrality below two. We exclude these from our regression when considering diversity metrics as independent variables. is the average pairwise distance between any two projects in : where and are projects in , and and are their vector representations from the Node2Vec model. The distance between and is the negation of the cosine similarity of the two vectors. This step resulted in computing the three variables listed in Table 2 under Diversity Variables (original).
3.5. Dimensionality Reduction
@@
Since they are designed to capture different concepts, we run PCA separately for the three degree variables and the three diversity variables. As standard, we first log-transformed all input variables and scaled them to a mean of zero and a standard deviation of one.
@@
The PCA resulted in three components for the degree variables and another three for the diversity variables, with the proportion of variance explained by each component listed in Table 1 (top). Inspecting the table, we observe that the first two PCs cumulatively explain over 80% of the variance in both groups, thus we decided to retain only two components each for the degree and diversity variables. That is, we will use these PCs (listed under Degree Variables (post-PCA) and Diversity Variables (post-PCA) in Table 2) instead of the original variables in our regression models below.
@@
To interpret the two PCs we turn to Table 1 (bottom), which lists the loadings of the principal components onto the original variables.444Principal components are linear combinations of the original variables, with the loadings indicating the contributing coefficients. For example, the first PC for the degree variables, , uses the coefficients in the first column in Table 1 (bottom): . Inspecting the table we make the following observations. First, the loadings onto the first principal component (PC1) are relatively consistent across networks for both the degree and the diversity variables. Thus, we interpret PC1 to represent the average degree (or diversity of ties) of the project across the three networks. For example, drawing on our theoretical framework (Section 2), a project for which is high is expected to have many ties (on average, across the three networks), i.e., its core developers should have more sources from which to draw inspiration. Similarly, the ties of a project for which is high are expected to span more of the knowledge embedding space we used to estimate tie diversity, i.e., its core developers could have access to more varied sources to draw inspiration from.
| Out-deg. centrality | Diversity index | ||||||
| PC1 | PC2 | PC3 | PC1 | PC2 | PC3 | ||
| Variance Explained | 0.64 | 0.22 | 0.14 | 0.51 | 0.29 | 0.19 | |
| Cumulative Variance | 0.64 | 0.86 | 1.00 | 0.51 | 0.81 | 1.00 | |
| 0.60 | 0.45 | 0.67 | 0.63 | 0.36 | 0.69 | ||
| 0.61 | 0.28 | 0.74 | 0.65 | 0.24 | 0.72 | ||
| 0.52 | 0.85 | 0.11 | 0.43 | 0.90 | 0.08 | ||
@@
Second, the loadings onto the second component (PC2) show a descending trend across networks, again similarly for both sets of variables. Notably, the highest loading on the degree (or diversity) comes from the star network, followed by the issue network, while the lowest comes from the commit network. Thus, we interpret PC2 to represent the strength of network ties in the degree (or diversity of ties) metric, i.e., the strength of weak ties. A project for which is high is expected to get more of its connectivity through the star network. Analogously, relatively more of the diversity of knowledge accessible to core developers in a project for which is high can be attributed to the star network (weak ties) compared to the commit or issue networks (stronger ties).
@@
It may seem counterintuitive to reason about PC1 and PC2 jointly, especially in a regression modeling framework. What does it mean to vary PC2 from low to high, for instance, while holding PC1 fixed? Can PC1 and PC2 even vary independently? They can in theory, because they are orthogonal by construction, but are the combinations of low-high values of PC1 and PC2 observable in the real-world data? After all, regression is an interpolation mechanism. We illustrate this space of interpretations of PC1 and PC2 for the diversity variables with the artificial examples in Figure 3 (see Supplementary for real-world examples in each quadrant). In the top left quadrant (PC1 and PC2 both high), the project core developers can tap into a diverse knowledge space via either strong or weak ties. In the top right (PC1 high and PC2 low), the overall diversity of information accessible is just as high, but it’s the strong ties that are the source of it. In the bottom left (PC1 low and PC2 high), the overall diversity of information available is lower, but whatever diversity there is, it can be attributed mostly to the weak ties. Finally, in the bottom right (PC1 and PC2 both low), weak ties contribute little to an otherwise low diversity of information.
@@
Going back to our hypotheses (Section 2), we expect the left quadrants, where weak ties are strong, to be most associated with innovativeness. As an anecdote, consider the example in Figure 1 on the first page. The GeoNotebook project is a Jupyter notebook-based environment for interactive visualization and analysis of geographic data. Interestingly, GeoNotebook’s core developers have previously starred (without otherwise contributing to) many GitHub projects spanning a variety of seemingly related topics, including a collection of free APIs, example data science notebooks, reproducible workflows for Jupyter notebooks, visualization tools, time series analysis tools, geographic data analysis tools, and infrastructure tools for interactive packet manipulation. One can expect that this diverse space of ideas provided some inspiration, knowingly or unknowingly, for the design and implementation of GeoNotebook. Our analysis below tests to what extent there is evidence supporting this mechanism at scale, beyond this anecdote.
@@
| Innovativeness | The extent to which a project imports packages that were typically not combined in the past, computed based on all the packages imported until the end of 2021. |
@@
| The total number of core developers in the project, as computed before the end of 2021. | |
| The number of software packages that the project imported before the end of 2021. | |
3.6. Project Innovativeness Measure
For each project, we extract the sequence of packages imported555We ignore built-in Python libraries like Fang et al. (2024) did, and consider only packages published in the PyPI registry, using the explicit links back to GitHub recorded there. (precomputed in World of Code) and use a skip-gram model to generate positional embeddings for each package based on these dependency relationships. These embeddings place the package in an optimal location relative to all other packages imported in the same project, based on their co-import relationships. Unlike the random walks used in the Node2Vec model to generate embeddings for the three interaction networks, there is no inherent order in the sequence of packages imported within the same project. Therefore, we set the window size in the skip-gram model to be sufficiently large, enabling it to consider all packages in the same sequence as its context when generating embeddings. Similar approaches to compute embeddings of software packages have been used before in software engineering research (Dey et al., 2021; Fang et al., 2023).
@@ 3.6. Project Innovativeness Measure
To estimate the atypicality of a combination of two packages, we compute the negation of their embeddings’ cosine similarity. Then, our measure of project innovativeness is the average pairwise atypicality of all pairs of packages imported by a project.
Validation. Although measuring project novelty in terms of the atypicality of package combinations therein has solid theoretical foundation (Section 2) in addition to application in past empirical studies of open-source software (Fang et al., 2024), this likely remains the most controversial part of our methodology. Therefore, we perform two validation checks, comparing to an existing dataset of “awesome” projects and to the previous measure by Fang et al. (2024).
“Awesome” lists are community-curated lists of software. Many exist for different programming languages, platforms, application domains, etc. One can propose new entries via a pull request (PR), and typically some minimum amount of support (+1s) is required for the PR to be merged. Being innovative isn’t formally a requirement for being “awesome,” although as part of the same PRs submitters usually argue that the proposed entries are novel (e.g., PRs to the Python list contain a section titled “What’s the difference between this Python project and similar ones?”)
We scraped software mentions from a popular aggregator of “awesome” lists (Sorhus, 2025) (350k+ GitHub stars at the time of writing), and identified 760 Python projects matching our sampling criteria (§3.2), and for which we could compute atypicality scores (e.g., having at least two import-based dependencies we could identify), corresponding to about 2% of our sample. Comparing the atypicality scores between the “awesome” and remaining projects in our sample (Figure 4), we observe a statistically significant difference in means (two-sample independent t-test ), with a non-trival effect size: a project with a median atypicality score in the non-selected group will only be among the bottom 35% ordered by atypicality score among the “awesome” projects. Further validation suggests this effect is not confounded by variables such as project size (Supplementary). Therefore, we conclude that our innovativeness measure reflects to some extent the developers’ perception of “awesome” projects, which usually includes a dimension of novelty.
In addition, we compute the correlation between the atypicality scores obtained using our skip-gram model and those generated using the original measure by Fang et al. (2024). Results show a moderate-to-strong (Schober et al., 2018) linear (Pearson) correlation of 0.65. Therefore, we conclude that our innovativeness measure correlates to that of prior work to a large degree. Furthermore, we test the robustness of our conclusions to variations in the innovativeness measure, and find that the estimated regression coefficients for the degree and diversity variables behave consistently with both measures. Thus, we only present the results obtained with our proposed measure below, but include the corresponding regression results with the original measure by Fang et al. (2024) in Supplementary.
3.7. Regression Modeling Considerations
@@
To test our three hypotheses, we use fixed-effects linear regression (one data point per project), where the response variable is our measure of innovativeness, and our main explanatory variables are the degree and diversity principal components discussed above: captures H (volume of interactions the core developers had with other projects in the previous year), captures H (diversity of information accessible), and captures H (strength of weak ties). The remaining principal component, , does not directly map to one of our hypotheses, but can help distinguish whether it’s the volume of weak ties, or their strength, that explains more variance in our outcome measure.
@@
We also control for several important confounding variables (Table 2): the repository owner’s social standing (), the project team size (), the complexity of the codebase in terms of number of packages imported (), whether the project is owned by an organizational account (Org_owned), and the project age (, modeled as a categorical variable). For model fit and diagnostics we followed standard practice, e.g., we first log-transformed the variables with skewed distributions to reduce heteroskedasticity (Kaufman, 2013); see replication package.
4. Results
@@
Next we present the main results from the regression analysis formally testing our hypotheses, summarized in Table 3, as well as a series of robustness checks.
@@
Main Regression Results. Model I (all projects) and Model II (projects with a minimum out-degree centrality of two across all three networks) shed light on the impact of degree variables. Model I reveals a significant positive effect from average degree centrality, whereas Model II shows a negative effect. Nevertheless, effect sizes for both models are relatively small. For instance, in Model I, transitioning a project’s from the 25th percentile (-1.09) to the 75th (0.53) results in a mere 0.002 increase in project innovativeness. Considering that the 25th percentile of a project’s innovativeness score is -0.33, and the 75th is -0.12 in our sample, the change in project innovativeness due to variations in average degree centrality is practically negligible. Therefore, H is not supported.
@@
In Model III, we observe a positive association between the average diversity () of ties originating from the focal project and project innovativeness. Moreover, this effect is notably stronger – nearly seven times greater – compared to that of average degree in Models I and II. Transitioning a project from the 25th percentile (-0.89) to the 75th (0.86) in terms of corresponds to an increase of approximately 0.012 in project innovativeness, representing roughly a 4% change in the distribution (e.g., a shift from the median project in terms of novelty to the 54th percentile). We conclude that H is supported.
@@
Additionally, we observe a significant effect for the variable. This suggests that, when controlling for the average diversity across the three networks, higher diversity in the “weak ties network” (i.e., star network), or stronger weak ties, corresponds to increased project innovativeness. In practical terms, the effect size for the variable is comparable, albeit somewhat smaller, to that of the variable above. We interpret this as supporting evidence for H.
@@
When incorporating both degree and diversity variables in the same regression model, as in Model IV, the positive effects of and remain significant. This suggests a consistent effect stemming from tie diversity and strength of weak ties. Nevertheless, we also note a significant negative effect for and after adjusting for project diversity. These negative effects suggest that, while holding tie diversity constant, higher out-degree centrality and more ties in the “weak network” are associated with decreased project innovativeness. This observation could be attributed to the increased constraints within a local community associated with higher out-degree, as discussed by Burt (Burt, 2000). Such constraints can lead to a reduced inclination for combining packages of diverse functions within the project, consequently contributing to the reduced novelty of the project. More research is needed to further test these effects.
@@
Robustness Checks. In addition to the regressions above, we test how robust our conclusions about the effects of and are to differences in operational decisions.
@@
Does Time of Measurement Matter? “No.” We segmented our dataset by project inception year and conducted regression analyses with the specification of Model III in Table 3 on projects initiated in each respective year. Figure 5(a) shows the resulting estimated coefficients for the two diversity variables. As illustrated in the graph, the positive relationship between the average diversity of knowledge access across the three networks (represented by the blue line) remains consistent over the years, with the average estimated coefficient being statistically significantly above zero in most years. The estimated coefficient for appears to also increase in recent years, though this difference is not statistically significant.
@@
The positive relationship between diversity from weak ties and project innovativeness is generally consistent (the red line in Figure 5(a)). The estimated coefficient is above zero in all ten years observed, and the statistically insignificant coefficient (i.e., the lower bound of the 95% confidence interval is below zero) is likely attributable to the limited number of projects in each annual cohort (i.e., fewer than 4,000 projects in each year before 2016).
@@
Does Project Size Matter? “No.” Similarly, we split projects into subsets based on the number of core developers and assessed the impact of the diversity variables in each cohort. As shown in Figure 5(b), the effect of remains consistent across projects of varying team sizes, confirming the robustness of our findings. It is noteworthy that the effect of becomes insignificant for projects with four or more core developers. This may be attributed to the limited number of project samples within these categories.
@@
Does Project Ownership Matter? “No.” Lastly, we re-estimated the effects of the diversity variables separately for projects owned by organizational and individual accounts. Our analysis (Figure 5(c)) shows that both average diversity across networks and diversity from weak ties have significantly positive effects for both categories. Furthermore, the difference in effect size between organization-owned and individual-owned projects is not statistically significant.
5. Discussion